a)     Suppose P (x, y) denotes the equation y=x+8, what will the truth values of the Propositions P (2,2), P (0,4).


1
Expert's answer
2021-12-15T10:05:07-0500

Suppose "P (x, y)" denotes the equation "y=x+8." Let us find the truth values of the Propositions "P (2,2), P (0,4)." Since "2\ne 10=2+8," we conclude that "P(2,2)" is false. Taking into account that "4\ne 8=0+8," we conclude that "P(0,4)" is also false.